# Enhanced sensing of Stark weak field under the influence of Aubry-André-Harper criticality

### v1

@@ I Introduction
QMB systems hosting distinct quantum phases can be used as a sensing probe. The unknown parameter to be estimated is encoded in the ground state or any excited state. These types of sensors may achieve HL at the criticality [39, 40, 41]. In order to estimate the parameter for these types of sensors we need to prepare a state and then adiabatically tune the parameter near criticality. This approach is referred as adiabatic sensing. In the critical region the system is extremely sensitive to even a small change of the parameter. Consequently, the system has drastic response to the parameter change. As for an example, due to a self-dual symmetry, the AAH model shows localization-delocalization transition at a finite strength of the potential amplitude. This transition characterized by a quasi-periodic potential can be used as a resource for achieving quantum-enhanced parameter sensing that attains HL [42]. Recently, there has been proposals for quantum sensors using stark localization. Interesingly, one can can achieve super Heisenberg limit via Stark localization transition [43]. The QMB sensors have various applications in quantum technology [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. At the critical point or transition point the fidelity susceptibility scales with system size $L$ as $L^{2/d\nu}$, where $d$ represents the spatial dimensionality of the system, and $\nu$ is the critical exponent associated with the localization length, $\zeta$, near criticality ($h_{c}$), such that $\zeta\sim|h-h_{c}|^{-\nu}$. At the transition point the precision of the unknown parameter $h$ scales with system size as $L^{2/\nu}$ [63, 64]. Experimentally, it is difficult to measure QFI. Therefore we need to find an experimentally measurable operator. For operators, one needs to compute operator based Fisher information (OFI) which is basically defined via signal-to-noise ratio. It can be shown that OFI is bounded by QFI [65]. In principle, the OFI saturates QFI when optimized over all possible operators.
@@
In the following, section I presents the introduction. Section II presents fidelity-based definition of QFI and OFI. The system is introduced in section III. Section IV demonstrates the behavior of QFI, OFIs, and their scaling with system size for a single-particle probe. Section V covers the QFI, OFIs, and their scaling for a many-body half-filled probe. Finally, the discussion is provided in section VI.
Figure 1: Single party QFI: (a) presents QFI against the stark field $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (b) shows the scaling plot. The maroon circles are QFI for different system sizes $L=21,34,55,89,144,233,377,610$ at $h=10^{-9}$, and the dotted line is the best fit $F_{Q}\sim L^{6.7}$. Both plots are shown in log-log scale. A configuration averaging over $\phi$ is performed with 8000 random samples.
@@ III System
where $\hat{c}_{i}^{\dagger}(\hat{c}_{i})$ is creation (annihilation) operator and $L$ denotes the system size. The first term in $\hat{H}_{0}$ is the kinetic energy term. The lattice is subjected to two different kinds of onsite potentials. The second term in $\hat{H}_{0}$ corresponds to a quasiperiodic onsite potential and $H_{1}$ has contribution from the onsite potential energy due to the linear gradient field with strength $h$. Here $\omega$ is taken as $\omega=F_{n}/F_{n+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. For $n\to\infty$ limit, $\omega=\lim_{n\to\infty}F_{n}/F_{n+1}\to(\sqrt{5}-1)/2$, which is golden ratio. In presence of the quasiperiodic potential, appropriate scaling emerges for the system sizes belonging to the fibonacci series [80]. $\phi$ is a phase which is chosen randomly from a uniform distribution in [0,1]. We consider open boundary condition (OBC). The Hamiltonian ($\hat{H}$) reduces to pure stark case for $V=0$. Then the system is localized in the zero field limit, i.e., $h\to 0$. At this transition point the fidelity susceptibility scales with system size as $\chi_{Q}\sim L^{2/\nu}$ with $\nu=0.33$ for single particle ground state [43]. Now, when the system is additionally at the criticality of AAH, i.e., $V=2$, the corresponding Hamiltonian is under the influence of two distinct onsite potentials. The scaling exponent associated with the localization-delocalization is reported as $\nu=0.29$ [80].
Figure 2: Single party OFI: (a) presents OFI, $F_{cdw}$, for observable $\hat{O}_{cdw}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangulat up), 987 (green triangular left). (b) presents OFI, $F_{H_{2}}$ for observable $\hat{O}_{H_{2}}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (urquoise diamond), 610 (black triangulat up), 987 (green triangular left). (c) illustrates the scaling of OFI corresponding to the operator $\hat{O}_{cdw}$, $F_{cdw}$ (circles) and $F_{H_{2}}$ (stars), respectively. The straight lines are the best fits. Whereas $F_{cdw}$ scales as $F_{cdw}\sim L^{6.2}$, $F_{H_{2}}$ saturates the scaling of the QFI, i.e., $F_{H_{2}}\sim L^{6.7}$.
@@ IV Single particle system as probe
IV.1 Quantum Fisher Information
We first consider the single-particle scenario. Our goal is to perform precise measurement of $h$ by setting the AAH potential at criticality. All eigen functions undergo delocalized to localized phase transition for this system. In the thermodynamic limit ($L\to\infty$) the transition point takes place at $h\to 0$ [43]. To address the question of whether the sensitivity of $h$ can be enhanced using the criticality of the AAH model, we investigate and calculate both the QFI and the OFI in details.
In Fig. 1(a) we show QFI as a function of the stark strength $h$ for $V=2$, when the ground state of the system is exploited as a probe. We study various system sizes $L=144,233,377,610,987$. Throughout the calculation, we have performed averaging over $\phi$ for our single particle probe with OBC, where $\phi$ is sampled randomly from a uniform distribution (see Eq. (III)). The figure illustrates that for a finite size system, as $h$ enhances adiabatically, the QFI remains initially flat indicating an extended nature of the wavefunctions. The QFI starts decreasing beyond this flat region, which indicates entrance of the system into the localized phase. So, in the extended regions corresponding to the finite systems, there is a scaling of QFI with $L$ whereas in the localized region, QFI becomes scale invariant, i.e., system size independent. Fig. 1(b) shows the scaling of QFI with finite system sizes in the extended region. The maroon circle is the value of QFI for different system sizes and the maroon dot-dot line is a fitting function $F_{Q}\sim L^{\beta}$ with $\beta=6.7$.
@@ IV.1 Quantum Fisher Information
IV.2 Operator Fisher Information
@@ VI Conclusions
This work proposes to use different classes of localization inducing potentials that can enhance the precision of parameter estimation even further. Our studies reveal that near the criticality of the AAH potential, better precision in estimating weak fields of Stark strength can be achieved with the QFI scaling as $L^{6.7}$ in comparison to the pure Stark system for which the QFI is reported to scale as $L^{6}$. In addition, this work proposes an experimentally accessible observable, the charge-density-wave (CDW) order denoted as $\hat{O}_{cdw}$. Remarkably, this observable also exhibits super-Heisenberg scaling with the QFI scaling as $F_{cdw}\sim L^{6.2}$. Moreover, another observable, $\hat{O}_{H_{2}}$, has been identified for which the associated OFI not only saturates to the QFI but also matches the scaling behavior of the QFI. For many-body systems, particularly under the condition of half-filling, the scaling of the QFI with system size is found as $F_{Q}\sim L^{6.6}$. Similarly, for the half-filled case, the OFI corresponding to the operator $\hat{O}_{H_{2}}$ exhibits the same scaling behavior as the QFI, that is $F_{H_{2}}\sim L^{6.6}$.
In summary, our work proposes an experimentally realizable protocol for achieving enhanced sensing of the Stark weak field by invoking another type of localization-inducing potential in form of the AAH potential and by exploiting the AAH criticality. One may, in principle, envision the underlying strength of the concept which can, in principle, be generalized for criticality based quantum sensors – In context of precision estimation of an unknown parameter within the QMB platform, it might be advantageous to introduce additional control parameters, that, in turn, introduces new kinds of quantum criticality, for obtaining enhanced quantum advantage by utilizing the competing effects of multiple critical points.

### v2

@@ I Introduction
QMB systems hosting distinct quantum phases can be used as a sensing probe. The unknown parameter to be estimated is encoded in the ground state or an excited state. These types of sensors may achieve HL at the criticality [39, 40, 41]. In order to estimate the parameter for these types of sensors we need to prepare a state and then adiabatically tune the parameter near criticality. This approach is referred as adiabatic sensing. In the critical region, the system is extremely sensitive to even a small change of the parameter. Consequently, the system has drastic response to the parameter change. As for an example, due to a self-dual symmetry, the AAH model shows localization-delocalization transition at a finite strength of the potential amplitude. This transition characterized by a quasi-periodic potential can be used as a resource for achieving quantum-enhanced parameter sensing that attains HL [42]. Recently, there has been proposals for quantum sensors using stark localization. Interesingly, one can can achieve super Heisenberg limit via Stark localization transition [43]. The QMB sensors have various applications in quantum technology [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62]. At the critical point or transition point the fidelity susceptibility scales with system size $L$ as $L^{2/d\nu}$, where $d$ represents the spatial dimensionality of the system, and $\nu$ is the critical exponent associated with the localization length, $\zeta$, near criticality ($h_{c}$), such that $\zeta\sim|h-h_{c}|^{-\nu}$. At the transition point the precision of the unknown parameter $h$ scales with system size as $L^{2/\nu}$ [63, 64]. Experimentally, it is difficult to measure QFI. Therefore we need to find an experimentally measurable operator. For operators, one needs to compute operator-based Fisher information (OFI) which is basically defined via signal-to-noise ratio. It can be shown that OFI is bounded by QFI [65]. In principle, the OFI saturates QFI when optimized over all possible operators.
@@
In the following, section I present the introduction. Section II presents the fidelity-based definition of QFI and OFI. The system is introduced in section III. Section IV demonstrates the behavior of OFIs and their scaling with system size for a single-particle probe. Section V covers the QFI, OFIs, and their scaling for a many-body half-filled probe. Finally, the discussion is provided in section VI.
@@ III System
where $\hat{c}_{i}^{\dagger}(\hat{c}_{i})$ is creation (annihilation) operator and $L$ denotes the system size. The first term in $\hat{H}_{0}$ is the kinetic energy term. The lattice is subjected to two different kinds of onsite potentials. The second term in $\hat{H}_{0}$ corresponds to a quasiperiodic onsite potential and $H_{1}$ has a contribution from the onsite potential energy due to the linear gradient field with strength $h$. Here $\omega$ is taken as $\omega=F_{n}/F_{n+1}$, where $F_{n}$ is the $n^{\text{th}}$ Fibonacci number. For $n\to\infty$ limit, $\omega=\lim_{n\to\infty}F_{n}/F_{n+1}\to(\sqrt{5}-1)/2$, which is golden ratio. In the presence of the quasiperiodic potential, appropriate scaling emerges for the system sizes belonging to the Fibonacci series [80]. $\phi$ is a phase which is chosen randomly from a uniform distribution in [0,1]. We consider open boundary condition (OBC). The Hamiltonian ($\hat{H}$) reduces to pure stark case for $V=0$. Then the system is localized in the zero field limit, i.e., $h\to 0$. At this transition point the fidelity susceptibility scales with system size as $\chi_{Q}\sim L^{2/\nu}$ with $\nu=0.33$ for single particle ground state [43]. Now, when the system is additionally at the criticality of AAH, i.e., $V=2$, the corresponding Hamiltonian is under the influence of two distinct onsite potentials. The scaling exponent associated with the localization-delocalization is reported as $\nu=0.29$ [80].
Figure 1: Single party OFI: (a) presents OFI, $F_{cdw}$, for observable $\hat{O}_{cdw}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (b) presents OFI, $F_{H_{2}}$ for observable $\hat{O}_{H_{2}}$ with respect to $h$ for system size L = 144 (blue circle), 233 (red square), 377 (turquoise diamond), 610 (black triangular up), 987 (green triangular left). (c) illustrates the scaling of OFI corresponding to the operator $\hat{O}_{cdw}$, $F_{cdw}$ (circles) and $F_{H_{2}}$ (stars), respectively. The straight lines are the best fits. Whereas $F_{cdw}$ scales as $F_{cdw}\sim L^{6.2}$, $F_{H_{2}}$ saturates the scaling of the QFI, i.e., $F_{H_{2}}\sim L^{6.7}$.
@@ IV Single particle system as probe
We first consider the single-particle scenario. We aim to precisely measure $h$ by setting the AAH potential at criticality. All eigenfunctions undergo delocalized to localized phase transition for this system. In the thermodynamic limit ($L\to\infty$) the transition point takes place at $h\to 0$ [43]. To address whether the sensitivity of $h$ can be enhanced using the criticality of the AAH model, we investigate and calculate the OFI in detail.
The scaling of QFI with finite system sizes at the extended region is $F_{Q}\sim L^{\beta}$ with $\beta=6.7$ which is shown in the reference [80]. In this section, we have shown that the experimentally measurable operator can also achieve super-Heisenberg scaling for measurement precision of the parameter $h$.
@@
IV.1 Operator Fisher Information
@@ VI Conclusions
This work proposes to use different classes of localization inducing potentials that can enhance the precision of parameter estimation even further. Near the criticality of the AAH potential, better precision in estimating weak fields of Stark strength can be achieved with the QFI scaling as $L^{6.7}$ in comparison to the pure Stark system for which the QFI is reported to scale as $L^{6}$. In addition, this work proposes an experimentally accessible observable, the charge-density-wave (CDW) order denoted as $\hat{O}_{cdw}$. Remarkably, this observable also exhibits super-Heisenberg scaling with the QFI scaling as $F_{cdw}\sim L^{6.2}$. Moreover, another observable, $\hat{O}_{H_{2}}$, has been identified for which the associated OFI not only saturates to the QFI but also matches the scaling behavior of the QFI. For many-body systems, particularly under the condition of half-filling, the scaling of the QFI with system size is found as $F_{Q}\sim L^{6.6}$. Similarly, for the half-filled case, the OFI corresponding to the operator $\hat{O}_{H_{2}}$ exhibits the same scaling behavior as the QFI, that is $F_{H_{2}}\sim L^{6.6}$.
In summary, our work proposes an experimentally realizable protocol for achieving enhanced sensing of the Stark weak field by invoking another type of localization-inducing potential in the form of the AAH potential and by exploiting the AAH criticality. One may, in principle, envision the underlying strength of the concept which can, in principle, be generalized for criticality-based quantum sensors – In the context of precision estimation of an unknown parameter within the QMB platform, it might be advantageous to introduce additional control parameters, that, in turn, introduces new kinds of quantum criticality, for obtaining enhanced quantum advantage by utilizing the competing effects of multiple critical points.